The Sum of three consecutive even numbers is 258? What are the numbers
Found 2 solutions by vasumathi, amoresroy:
Answer by vasumathi(46)

Solution : The three consecutive even numbers in terms of x are
(x-2) , x ,(x+2)
their sum is...
x-2 + x + x+2 = 3x
It is given that their sum is 258
so..
3x = 258
x = 258/3= 86
so x-2 = 86-2 = 84
and x+2 = 86+2 = 88
Then the three consecutive even numbers are
84,86,88

Answer by amoresroy(361)

Let x = the 1st even number
x + 2 = the 2nd even number
x + 4 = the 3rd evn number
x + x + 2 + x + 4 = 258
3x = 258 - 6
3x = 252
x = 84
The numbers are 84,86, & 88